A bicycle leaves town at 10 mph at 8 am. A car leaves town going the same direction at 11am going 50 mph. How many minutes after 11 am will it be when the car overtakes the bicycle?
Answer by checkley77(12844)

D=RT
D=10T FOR THE BICYCLE.
D=50(T-3) FOR THE CAR.
10T=50(T-3)
10T=50T-150
10T-50T=-150
-40T=-150
T=-150/-40
T=15/4=3.75 TRAVEL TIME FOR THE BICYCLE.
3.75-3=.75 TRAVEL TIME FOR THE CAR.
PROOF:
10*3.75=50*.75
37.5=37.5